# Reflected multi-entropy and its holographic dual

### v1

@@ Abstract
We introduce a mixed-state generalization of the multi-entropy through the canonical purification, which we called reflected multi-entropy. We propose the holographic dual of this measure. For the tripartite case, a field-theoretical calculation is performed using a six-point function of twist operators at large $c$ limit. At both zero and finite temperature, the field-theoretical results match the holographic results exactly, supporting our holographic conjecture of this new measure.
I Introduction
@@
Quantum entanglement [1, 2] has recently emerged as a key tool in understanding the AdS/CFT correspondence [3, 4, 5], which remains one of our most successful frameworks for probing quantum gravity. Remarkably, the Ryu-Takayanagi formula [6, 7, 8, 9] provides a means to calculate entanglement entropy in conformal field theories (CFTs) via the minimal area of a codimension-2 surface in anti-de Sitter (AdS) space, establishing a precise link between spacetime geometry and quantum entanglement. However, entanglement entropy exclusively captures quantum entanglement for pure states. To extend this connection to mixed states, holographic duals for several other measures, such as the entanglement of purification [10, 11, 12], reflected entropy [13, 14, 15], logarithmic negativity [16, 17, 18, 19, 20], and also odd entanglement entropy [21, 22], have recently been studied. Another important direction lies in the investigation of multipartite correlation measures [23, 24, 25, 26, 27, 28, 29, 30] and their holographic duals play a critical role in understanding the emergence of bulk spacetime from many-body quantum entanglement on the boundary. The aim of this letter is to bridge the gap between multipartite correlations in mixed states and spacetime geometry. The multipartite entanglement of purification has been investigated in [31, 32] and the multipartite generalization of reflected entropy has been studied in [33, 34].
@@
For a given multipartite mixed state, we can canonically purify it to obtain a pure state [13]. Using the recent proposal for pure state multipartite entanglement measure [35] based on the replica trick and permutation invariance, one can compute the so-called multi-entropy for the purified state. We propose this as an intrinsic multipartite measure for the original mixed state, which we term reflected multi-entropy. Unlike multi-entropy [36, 37, 38, 39], reflected multi-entropy is intrinsically well-defined in the ultraviolet, thus circumventing the divergence issue that plagues entanglement entropy as well as multi-entropy in QFT. In the remainder of this letter, we first introduce reflected multi-entropy as the mixed-state generalization of multi-entropy, and then propose its holographic dual. For the tripartite case, we perform a field-theoretical calculation using a six-point function of twist operators and find that it matches the holographic result exactly, both at zero and finite temperature.
II A Generalization of multi-entropy
II.1 Review of multi-entropy
Let us recall some basic notions of entanglement entropy (EE) and Rényi entropy. Given a pure state |$\ket{\psi}_{AB}$ that describes two systems $A$ and $B$, the EE between $A$ and $B$ is defined as
@@ II.1 Review of multi-entropy
where $\rho_{A}=\text{Tr}_{B}\ket{\psi}\bra{\psi}_{AB}$ is the reduced density matrix. Rényi entropy is a one-parameter generalization of EE
@@
with $\sigma=(123\cdots n)$ the permutation acting on the index of replicas. Guided by this key observation, in [35] the authors defined the multi-entropy $S^{(\mathtt{q})}$, which generalizes the concept of EE to $\mathtt{q}$-partite cases through the permutation and contraction of indices. Let us focus on tripartite pure state |$\ket{\psi}_{ABC}$ to illustrate.
Like Rényi entropy, the entanglement measure $S_{n}^{(\mathtt{q}=3)}$ should be determined by the choice of permutation $(\sigma_{A},\sigma_{B},\sigma_{C})$ acting on the indices of different replicas, i.e., the way of contracting multiple density matrices. Note that there is a equivalence relation
$(\sigma_{A},\sigma_{B},\sigma_{C})\sim(\sigma_{A},\sigma_{B},\sigma_{C})\cdot g$ (4)
allowing us to consider only $(\mathtt{q}-1)$-parties, i.e., we can always choose $g=\sigma_{A}^{-1}$ to make $\sigma_{A}=\text{id}$ and the entanglement measure is only determined by $(\sigma_{B},\sigma_{C})$. [^1] In this letter, we treat all the subsystems symmetrically. Both $B$ and $C$ have $n$ replicas and $(\sigma_{B},\sigma_{C})$ are chosen to be cyclic permutations of order $n$. In total, we have $n^{\mathtt{q}-1}=n^{2}$ replicas, which form a $n\times n$ square lattice. The multi-entropy is defined by choosing the replica symmetry as $\mathbb{Z}_{n}\otimes\mathbb{Z}_{n}$ with the first/second $\mathbb{Z}_{n}$ acting on the subregion $B$/$C$ and cyclically permuting the replicas located in the same row/column of the lattice.
For general $n$ and $\mathtt{q}$-partite entanglement, the Rényi multi-entropy is defined as [^2] Compared with [35], we use a slightly different notation here. We denote the partition function on the $n^{\mathtt{q-1}}$-sheet Riemann surface by $\mathcal{Z}^{(\mathtt{q})}_{n^{\mathtt{q}-1}}$ instead of $\mathcal{Z}^{(\mathtt{q})}_{n}$ in [35].
$S^{(\mathtt{q})}_{n}=\frac{1}{1-n}\log(\mathcal{Z}^{(\mathtt{q})}_{n^{\mathtt{q}-1}}/(\mathcal{Z}_{1}^{(\mathtt{q})})^{n^{\mathtt{q}-1}})\ ,$ (5)
where $\mathcal{Z}^{(\mathtt{q})}_{n^{\mathtt{q}-1}}$ is the partition function on the $n^{\mathtt{q}-1}$-sheet Riemann surface with replica symmetry $\mathbb{Z}_{n}^{\otimes(\mathtt{q}-1)}$ as described above. The multi-entropy is defined by taking the $n\to 1$ limit
$S^{(\mathtt{q})}=\lim_{n\to 1}S^{(\mathtt{q})}_{n}\ .$ (6)
For $\mathtt{q}=3$ and $n=2$, the partition function contains 4 replicas with the gluing $\sigma_{B}=(13)(24)$ and $\sigma_{C}=(12)(34)$ as illustrated in FIG. 1 via the Euclidean path integral [40, 41, 42].
@@
This results in a contraction of four density matrices (here we abbreviate $\rho_{BC}$ as $\rho$)
$\mathcal{Z}^{(3)}_{2^{3-1}}=\rho_{\beta_{1}\chi_{1}}^{\beta_{3}\chi_{2}}\rho_{\beta_{2}\chi_{2}}^{\beta_{4}\chi_{1}}\rho_{\beta_{3}\chi_{3}}^{\beta_{1}\chi_{2}}\rho_{\beta_{4}\chi_{4}}^{\beta_{2}\chi_{3}}\ .$ (7)
The corresponding Rényi multi-entropy is given by
$S^{(3)}_{2}=\frac{1}{1-2}\log\left(\mathcal{Z}^{(3)}_{2^{3-1}}/(\mathcal{Z}_{1}^{(3)})^{2^{3-1}}\right)\ ,$ (8)
with $\mathcal{Z}_{1}^{(3)}=\rho^{\beta\chi}_{\beta\chi}$ serving as the normalization factor.
II.2 Mixed-state generalization of multi-entropy through canonical purification
Figure 2: Canonical purification of $\rho_{ABC}$. (a) Global pure state $\psi_{AcBaCb}$ defined on a circle. Tracing out $abc$ gives a mixed state $\rho_{ABC}$. (b) Picking up another copy of $\rho_{ABC}$ and gluing these two copies along $abc$ gives a big pure state |$\ket{\sqrt{\rho_{ABC}}}$ (9), which is the canonical purification of $\rho_{ABC}$.
A multipartite mixed state can be obtained by tracing out some part from a pure state. Again we will focus on tripartite case with $\mathtt{q}=3$ but it can be generalized to larger $\mathtt{q}$ straightforwardly. Starting from a pure state $\psi_{AcBaCb}\in\mathcal{H}_{AcBaCb}$ and tracing out $abc$, we obtain a mixed state $\rho_{ABC}$, which can be canonically purified by doubling the Hilbert space (see FIG. 2 for a schematic diagram of this purification process). For more discussions on canonical purification we refer to [13]. The purified state can be expressed as
|$\begin{split}\ket{\sqrt{\rho_{ABC}}}&=\ket{\sqrt{\text{Tr}_{abc}\ket{\psi_{AcBaCb}}\bra{\psi_{AcBaCb}}}}\\ &\in\left(\mathcal{H}_{A}\otimes\mathcal{H}_{A}^{*}\right)\otimes\left(\mathcal{H}_{B}\otimes\mathcal{H}_{B}^{*}\right)\otimes\left(\mathcal{H}_{C}\otimes\mathcal{H}_{C}^{*}\right)\ .\end{split}$ (9)
@@ II.2 Mixed-state generalization of multi-entropy through canonical purification
$S_{R}^{(\mathtt{q}=3)}(A;B;C)=S^{(\mathtt{q}=3)}(AA^{*};BB^{*};CC^{*})_{\ket{\sqrt{\rho_{ABC}}}}\ ,$ (10)
with $S^{(\mathtt{q})}$ the original multi-entropy (6). The subscript $R$ of $S^{(\mathtt{q})}_{R}$ denotes “reflected”. [^3] The marriage with holographic purification was also mentioned in [36] without explicit construction.
@@
  1. (i)
    It reduces to twice the multi-entropy for pure state $\rho_{AB\cdots}=\ket{\psi}\bra{\psi}_{AB\cdots}$,
@@
  2. (ii)
    It reduces to the reflected entropy [13] for bipartite system $\rho_{AB}$,
@@
  3. (iii)
    For factorized density matrix $\rho_{AB\cdots}=\rho_{A}\otimes\rho_{B}\otimes\cdots$, $S_{R}^{(\mathtt{q})}(A;B;\cdots)$ vanishes,
@@
  4. (iv)
    $S_{R}^{(\mathtt{q})}(A;B;\cdots)$ is bounded from below by multi-entropy of purification,
[…]
@@
  5. (v)
    $S_{R}^{(\mathtt{q})}(A;B;\cdots)$ is bounded from below by half the multipartite reflected entropy [34] for holographic states,
[…]
III Holography of reflected multi-entropy
In the previous section we introduced the reflected multi-entropy as a multipartite measure for mixed states. Based on the canonical purification for holographic states [13] and the holographic proposal for multi-entropy [35], now we want to find the holographic dual for reflected multi-entropy. Consider a global pure state $\psi_{AcBaCb}$ defined on a circle as show in FIG. 2(a). This pure state has a classical bulk solution as its holographic dual. Now we trace out $abc$, which corresponds to retaining only the entanglement wedge for $\rho_{ABC}$ and prescinding other portion of the bulk. To perform the canonical purification, we pick up another copy of the entanglement wedge and glue these two copies along $\Gamma_{ABC}$, the Ryu-Takayanagi surface (RT surface) for $\rho_{ABC}$. The resulting geometry is drawn in FIG. 3(a).
Figure 3: (a) Canonical purification of $\rho_{ABC}$ and its holographic dual. Tracing out $abc$ corresponds to gluing along $\Gamma_{ABC}$, the RT surface for $\rho_{ABC}$ (blue curves). The orange web $\mathcal{W}$ is the minimal web in the bulk as the holographic dual of $S^{(3)}(AA^{*};BB^{*};CC^{*})$. (b) The single copy picture. The shadow area is the entanglement wedge of $A\cup B\cup C$ and the orange web is the holographic dual of $S_{R}^{(3)}(A;B;C)$.
To figure out the geometry dual of the reflected multi-entropy defined in subsection II.2, which now becomes the multi-entropy between $AA^{*}$, $BB^{*}$ and $CC^{*}$, we need to generalize the holographic proposal for multi-entropy [35] to multi-boundary pure states. Recall that the original holographic proposal for multi-entropy is a minimal bulk web $\mathcal{W}$ consist of minimal codimension-2 surfaces. For a single boundary pure state, $\mathcal{W}$ is anchored at the boundaries of all subsystems $A,B,\cdots$. This condition is relaxed in the case of multi-boundary, much like the RT surfaces for multi-boundary states. Another important condition is that $\mathcal{W}$ should contain sub-webs that are homologous to all the subsystems $A,B,\cdots$. This can be easily satisfied. It is then natural to propose the holographic dual of the reflected multi-entropy as the minimal surface web shown in FIG. 3(a). The situation is similar for the four-partite case as drawn in FIG. 4, but now there are two channels and one should pick the minimum of them.
Figure 4: Four-partite case.
More precisely, the holographic dual of the reflected multi-entropy $S_{R}^{(\mathtt{q})}(A;B;\cdots)$ is the minimal surface web $\mathcal{W}$ satisfying the following topological conditions:
@@ III Holography of reflected multi-entropy
    $\mathcal{W}$ is anchored at $\Gamma_{AB\cdots}$ which is the RT surface of the system $A\cup B\cup\cdots$.
@@
    $\mathcal{W}$ contains sub-webs that are homologous to all the subsystems $A,B,\cdots$.
@@
IV Computation in $\text{AdS}_{3}/\text{CFT}_{2}$
IV.1 Zero temperature case
@@
$\mathrm{d}s^{2}=\frac{\mathrm{d}x^{2}+\mathrm{d}y^{2}}{y^{2}}\ ,\quad x\in(-\infty,+\infty)\ \&\ y\in[0,+\infty)\ .$ (17)
We consider the reflected multi-entropy $S^{(3)}_{R}(A;B;C)$ between three subsystems $A,B,C$. $\rho_{ABC}$ is a mixed state obtained by tracing out $abc$ from the pure state $\psi_{AcBaCb}$ defined on $\mathbb{R}$ as shown in FIG. 5. The subsystems $A,B,C$ are chosen to be
$A=[x_{1},x_{2}]\ ,\quad B=[x_{3},x_{4}]\ ,\quad C=[x_{5},x_{6}]\ ,$ (18)
with $x_{i=1,\dots,6}$ parameterized by $\xi=a,b,c$
$\begin{split}x_{6}&=-x_{1}=b\ ,\\ x_{5}&=-x_{2}=a+r\ ,\\ x_{4}&=-x_{3}=a-r\ .\end{split}$ (19)
In the following we will calculate $S^{(3)}_{R}(A;B;C)$ using both holographic method and field theoretical method and find the precise agreement.
IV.1.1 Holographic calculation
As we have seen in section III, the holographic dual of $S^{(3)}_{R}(A;B;C)$ is given by the minimal web denoted by orange curves in FIG. 5. Due to the $\mathbb{Z}_{2}$-symmetry about the $x=0$ axis in FIG. 5, the junction $I$ is expected to be on $x=0$ and we denote it by $I(0,y_{I})$. For general two points $(x_{1},y_{1})$ and $(x_{2},y_{2})$ on the Poincaré half plane, the distance between them is
@@ IV.1.1 Holographic calculation
Therefore, the distance between $I$ and a point $(-a+r\cos\phi,r\sin\phi)$ on the RT surface of $c$ is
$D\bigl{(}(0,y_{I}),(-a+r\cos\phi,r\sin\phi)\bigr{)}\ ,$ (21)
which should be minimized over $\phi$. Similar distances between $I$ and the RT surface of $a,b$ can be worked out and we finally end up with the length of the bulk dual web
@@
where the first/second term in []delimited-[][$\cdots$][ ⋯ ] comes from the distance between $I$ and the RT surface of $b$/$\{a,c\}$. In (22) we did not minimize over $y_{I}$ and we will do that numerically when comparing with the field theory results (see the blue curves in FIG. 6).
IV.1.2 Field theoretical calculation
In this subsection we calculate $S_{R}^{(3)}(A;B;C)$ in conformal field theory. Using the replica trick, we first represent the reflected multi-entropy as a path integral on a specific Riemann surface. The example for $\{\mathtt{q}=3,n=3,m=2\}$ is shown in FIG. 10. Here a new replica number $m$ is introduced for canonical purification (see [13] and Appendix A for details).
Figure 5: The holographic dual of $S^{(\mathtt{q}=3)}_{R}(A;B;C)$.
The $n$-th Rényi version of reflected multi-entropy is given by
$S_{R}^{(\mathtt{q})}[n]=\frac{1}{1-n}\log\frac{\mathcal{Z}_{n^{\mathtt{q}-1},m}}{(\mathcal{Z}_{1,m})^{n^{\mathtt{q}-1}}}\ ,$ (23)
and the reflected multi-entropy can be obtained by taking the limit $n\to 1$ and $m\to 1$
$S_{R}^{(\mathtt{q})}=\lim_{m\to 1}\lim_{n\to 1}\frac{1}{1-n}\log\frac{\mathcal{Z}_{n^{\mathtt{q}-1},m}}{(\mathcal{Z}_{1,m})^{n^{\mathtt{q}-1}}}\ .$ (24)
Now we use the twist operator correlation function to express $\mathcal{Z}_{n^{\mathtt{q}-1},m}$, the path integral on a glued Euclidean spacetime with $mn^{\mathtt{q}-1}$ copies. The conformal weights $h_{i}$ of twist operators $\sigma_{i}(x_{i})$ are given by (see Appendix A for the derivation)
@@ IV.1.2 Field theoretical calculation
Figure 6: Comparison of the partial derivatives of $\frac{c}{3}L[\mathcal{W}]$ and $S_{R}^{(\mathtt{q}=3)}$ (divided by $c/3$) with respect to $\xi=a,b,r$. We take $a=20$, $b=100$ and $r$ ranging from $1$ to $10$.
When $\mathtt{q}=2$, one can check that (25) reduces to the conformal weight of the twist operators for the reflected entropy as we expected. [^4] For reflected entropy, we have [13] $h_{g_{A}}=h_{g_{A}^{-1}}=\frac{cn(m^{2}-1)}{24m}\ ,$ where $g_{A}$ and $g_{A}^{-1}$ are the twist operators located at the two endpoints of interval $A$. The conformal weight $h_{a}$ of the leading operators $\sigma_{a}$ in OPE $\sigma_{i}(x_{i})\sigma_{j}(x_{j})\to\sigma_{a=ij}(x_{a})$ are (see Appendix A for the derivation)
$h_{a=ij}=\frac{c}{12}\left(n^{\mathtt{q}-1}-n^{\mathtt{q}-3}\right)\ .$ (26)
The conformal weight (26) is twice the conformal weight of the twist operators for the original multi-entropy as we expected. [^5] The conformal dimension of the twist operators of multi-entropy is given in eq(29) of [38]. It can also be checked that taking $\mathtt{q}=2$, (26) reduces to twice of the conformal weight of the twist operators of EE.
@@
with $\sigma_{i}$ located at $x_{i}$. At large $c$ limit, the calculation of the six-point function (27) can be converted to solving a monodromy problem [43, 44], i.e., solving a second-order differential equation (46) with specific boundary conditions (49). Following the approach in [45] we numerically compute the six-point function (27) at large $c$ limit (the details are included in Appendix B) and the partial derivatives of $S_{R}^{(3)}(A;B;C)$ with respect to $a,b,r$ are plotted in FIG. 6.

### v2

@@ Abstract
We introduce a mixed-state generalization of the multi-entropy through the canonical purification, which we call reflected multi-entropy. We propose the holographic dual of this measure. For the tripartite case, a field-theoretical calculation is performed using a six-point function of twist operators at large $c$ limit. At both zero and finite temperature, the field-theoretical results match the holographic results, supporting our holographic conjecture of this new measure.
.1 Introduction
@@
Quantum entanglement [1, 2] has recently emerged as a key tool in understanding the AdS/CFT correspondence [3, 4, 5], which remains one of our most successful frameworks for probing quantum gravity. Remarkably, the Ryu-Takayanagi formula [6, 7, 8, 9] provides a means to calculate entanglement entropy (EE) in conformal field theories (CFTs) via the minimal area of a codimension-2 surface in anti-de Sitter (AdS) space, establishing a precise link between spacetime geometry and quantum entanglement. However, EE exclusively captures quantum entanglement for pure states. To extend this connection to mixed states, holographic duals for several other measures, such as the entanglement of purification [10, 11, 12], reflected entropy [13, 14, 15], logarithmic negativity [16, 17, 18, 19, 20], and odd entanglement entropy [21, 22], have recently been studied. Another important direction lies in the investigation of multipartite correlation measures [23, 24, 25, 26, 27, 28, 29, 30], whose holographic duals play a critical role in understanding the emergence of bulk spacetime from the boundary many-body quantum entanglement. The aim of this Letter is to bridge the gap between multipartite correlations in mixed states and spacetime geometry.
@@
For a given multipartite mixed state, we can canonically purify it to obtain a pure state [13]. Using the recent proposal for pure state multipartite entanglement measure [31] based on the replica trick and permutation invariance, one can compute the so-called multi-entropy for the purified state. We propose this as an intrinsic multipartite measure for the original mixed state, which we term reflected multi-entropy. Unlike multi-entropy [32, 33, 34, 35], reflected multi-entropy is intrinsically well-defined in the ultraviolet, thus circumventing the divergence issue that plagues EE as well as multi-entropy in quantum field theory. We note that while the multipartite generalizations of reflected entropy have been studied in [36, 37] [^1] A related measure called the multipartite entanglement of purification has been investigated in [39, 40]. , these attempts are fundamentally distinct from our approach. Crucially, unlike these previous measures in [36, 37]—which are defined through the von Neumann entropy on replica spaces—our proposed reflected multi-entropy is fundamentally rooted in multi-entropy, an intrinsic multipartite entanglement measure.
In the remainder of this Letter, we first introduce reflected multi-entropy as the mixed-state generalization of multi-entropy, and then propose its holographic dual. For the tripartite case, we perform a field-theoretical calculation using a six-point function of twist operators and find that it matches the holographic result, both at zero and finite temperature.
.2 A Generalization of multi-entropy
[…]
@@
[…]
@@
[…]
Like Rényi entropy, the entanglement measure $S_{n}^{(\mathtt{q}=3)}$ should be determined by the choice of permutation $(\sigma_{A},\sigma_{B},\sigma_{C})$ acting on the replica indices, i.e., the way of contracting multiple density matrices. Note that there is a equivalence relation $(\sigma_{A},\sigma_{B},\sigma_{C})\sim(\sigma_{A},\sigma_{B},\sigma_{C})\cdot g$ allowing us to set $\sigma_{A}=\text{id}$ by choosing $g=\sigma_{A}^{-1}$, reducing the problem to $(\mathtt{q}-1)$-parties with the entanglement measure depending solely on $(\sigma_{B},\sigma_{C})$  [^2] In this Letter, we treat all the subsystems symmetrically.. Both $B$ and $C$ have $n$ replicas and $(\sigma_{B},\sigma_{C})$ are chosen to be cyclic permutations of order $n$. In total, we have $n^{\mathtt{q}-1}=n^{2}$ replicas, which form a $n\times n$ square lattice. The multi-entropy is defined by choosing the replica symmetry as $\mathbb{Z}_{n}\otimes\mathbb{Z}_{n}$ with the first or second $\mathbb{Z}_{n}$ acting on the subregion $B$ or $C$ and cyclically permuting the replicas located in the same row or column of the lattice.
For general $n$ and $\mathtt{q}$-partite entanglement, the Rényi multi-entropy is defined as [31, 34]  [^3] Here we follow the definition in [34]. Compared with [31], we use a slightly different notation for the partition function on the replica space. We denote the partition function on the $n^{\mathtt{q-1}}$-sheet Riemann surface by $\mathcal{Z}^{(\mathtt{q})}_{n^{\mathtt{q}-1}}$ instead of $\mathcal{Z}^{(\mathtt{q})}_{n}$ in [31].
$S^{(\mathtt{q})}_{n}=\frac{1}{1-n}\frac{1}{n^{\mathtt{q}-2}}\log(\mathcal{Z}^{(\mathtt{q})}_{n^{\mathtt{q}-1}}/(\mathcal{Z}_{1}^{(\mathtt{q})})^{n^{\mathtt{q}-1}})\ ,$ (4)
where $\mathcal{Z}^{(\mathtt{q})}_{n^{\mathtt{q}-1}}$ is the partition function on the $n^{\mathtt{q}-1}$-sheet Riemann surface with replica symmetry $\mathbb{Z}_{n}^{\otimes(\mathtt{q}-1)}$ as described above. The multi-entropy is defined by taking the $n\to 1$ limit,
$S^{(\mathtt{q})}=\lim_{n\to 1}S^{(\mathtt{q})}_{n}\$. (5)
For $\mathtt{q}=3$ and $n=2$, the partition function contains four replicas with the gluing $\sigma_{B}=(13)(24)$ and $\sigma_{C}=(12)(34)$ as illustrated in Fig. 1 via the Euclidean path integral [43, 44, 45].
@@
This results in a contraction of four density matrices (here we abbreviate $\rho_{BC}$ as $\rho$),
$\mathcal{Z}^{(3)}_{2^{3-1}}=\rho_{\beta_{1}\chi_{1}}^{\beta_{3}\chi_{2}}\rho_{\beta_{2}\chi_{2}}^{\beta_{4}\chi_{1}}\rho_{\beta_{3}\chi_{3}}^{\beta_{1}\chi_{4}}\rho_{\beta_{4}\chi_{4}}^{\beta_{2}\chi_{3}}\ .$ (6)
The corresponding Rényi multi-entropy is then given by (4) with $\mathcal{Z}_{1}^{(3)}=\rho^{\beta\chi}_{\beta\chi}$ serving as the normalization factor.
[…]
Mixed-state generalization of multi-entropy through canonical purification. A multipartite mixed state can be obtained by tracing out some part from a pure state. Again we will focus on tripartite case with $\mathtt{q}=3$ but it can be generalized to larger $\mathtt{q}$ straightforwardly. Starting from a pure state $\psi_{AcBaCb}\in\mathcal{H}_{AcBaCb}$ and tracing out abcabc, we obtain a mixed state $\rho_{ABC}$, which can be canonically purified by doubling the Hilbert space (see Fig. 2 for a schematic diagram of this purification process). For more discussions on canonical purification we refer to [13]. The purified state can be expressed as
[…]
@@
[…]
with $S^{(\mathtt{q})}$ the original multi-entropy (5). The subscript $R$ in $S^{(\mathtt{q})}_{R}$ denotes “reflected”  [^4] The marriage with holographic purification was also mentioned in [32] without explicit construction..
@@
  1. (A)
@@
    $\text{pure state:}\ S_{R}^{(\mathtt{q})}(A;B;\cdots)=2S^{(\mathtt{q})}(A;B;\cdots)\ .$ (9)
  2. (B)
    It reduces to the reflected entropy [13] for bipartite system $\rho_{AB}$, i.e., $S_{R}^{(\mathtt{q}=2)}(A;B)=S_{R}(A;B)$.
  3. (C)
    It vanishes for factorized density matrix $\rho_{AB\cdots}=\rho_{A}\otimes\rho_{B}\otimes\cdots$,
    $\text{product state:}\ S_{R}^{(\mathtt{q})}(A;B;\cdots)=0\ .$ (10)
    However, $S_{R}^{(\mathtt{q})}(A;B;\cdots)$ can be nonzero for separable states. Consider a simple separable state:
@@
  4. (D)
    It is bounded from below by multi-entropy of purification,
    $\text{lower bound:}\ S_{R}^{(\mathtt{q})}(A;B;\cdots)\geq M_{P}(A;B;\cdots)\ ,$ (12)
@@
  5. (E)
    It is bounded from below by half the multipartite reflected entropy [37] for holographic states,
    $\text{lower bound:}\ S_{R}^{(\mathtt{q})}(A;B;\cdots)\geq\frac{1}{2}\Delta_{R}(A;B;\cdots)\ .$ (13)
.3 Holography of reflected multi-entropy
Based on the canonical purification for holographic states [13] and the holographic proposal for multi-entropy [31], now we want to find the holographic dual for reflected multi-entropy. Consider a global pure state $\psi_{AcBaCb}$ defined on a circle as shown in Fig. 2(a). This pure state has a classical bulk solution as its holographic dual. Now we trace out abcabc, which corresponds to retaining only the entanglement wedge for $\rho_{ABC}$ and prescinding the other portion of the bulk. To perform the canonical purification, we pick up another copy of the entanglement wedge and glue these two copies along the Ryu-Takayanagi surface (RT surface) for $\rho_{ABC}$. The resulting geometry is drawn in Fig. 3(a).
Figure 3: (a) Canonical purification of $\rho_{ABC}$ and its holographic dual. Tracing out abcabc corresponds to gluing along the RT surface for $\rho_{ABC}$ (blue curves). The orange web $\mathcal{W}$ is the minimal web in the bulk as the holographic dual of $S^{(3)}(AA^{*};BB^{*};CC^{*})$. (b) The single copy picture. The shadow area is the entanglement wedge of $A\cup B\cup C$ and the orange web is the holographic dual of $S_{R}^{(3)}(A;B;C)$.
To figure out the geometry dual of the reflected multi-entropy, which now becomes the multi-entropy between $AA^{*}$, $BB^{*}$, and $CC^{*}$, we need to generalize the holographic proposal for multi-entropy [31] to multi-boundary pure states. Recall that the original holographic proposal for multi-entropy is a minimal bulk web $\mathcal{W}$ consisting of minimal codimension-2 surfaces. For a single boundary pure state, $\mathcal{W}$ is anchored at the boundaries of all subsystems $A,B,\cdots$. This condition is relaxed in the case of multi-boundary, much like the RT surfaces for multi-boundary states. Another important condition is that $\mathcal{W}$ should contain subwebs that are homologous to all the subsystems $A,B,\cdots$. This can be easily satisfied. It is then natural to propose the holographic dual of the reflected multi-entropy as the minimal surface web shown in Fig. 3(a). More precisely, the holographic dual of the reflected multi-entropy $S_{R}^{(\mathtt{q})}(A;B;\cdots)$ is the minimal surface web $\mathcal{W}$ satisfying the following topological conditions:
@@
    $\mathcal{W}$ is anchored at the RT surface for $A\cup B\cup\cdots$.
@@
    $\mathcal{W}$ contains subwebs that are homologous to all the subsystems $A,B,\cdots$.
@@
.4 Computation in $\text{AdS}_{3}/\text{CFT}_{2}$
@@
$\mathrm{d}s^{2}=\frac{\mathrm{d}x^{2}+\mathrm{d}y^{2}}{y^{2}}\ ,\quad x\in\mathbb{R}\ \&\ y\in\mathbb{R}_{+}\ .$ (15)
We consider the reflected multi-entropy $S^{(3)}_{R}(A;B;C)$ between subsystems $A,B,C$. $\rho_{ABC}$ is a mixed state obtained by tracing out abcabc from the pure state $\psi_{AcBaCb}$ defined on $\mathbb{R}$ as shown in Fig. 4. The subsystems $A,B,C$ are chosen to be $A=[x_{1},x_{2}]$, $B=[x_{3},x_{4}]$, $C=[x_{5},x_{6}]$ with $x_{i=1,\dots,6}$ parametrized by $\xi=a,b,c$,
$\begin{split}x_{6}&=-x_{1}=b\ ,\\ x_{5}&=-x_{2}=a+r\ ,\quad x_{4}=-x_{3}=a-r\ .\end{split}$ (16)
In the following we will calculate $S^{(3)}_{R}(A;B;C)$ using both holographic method and field-theoretical method and find the agreement between them.
Figure 4: The holographic dual of $S^{(\mathtt{q}=3)}_{R}(A;B;C)$.
Holographic calculation. The holographic dual of $S^{(3)}_{R}(A;B;C)$ is the minimal web denoted by the orange curves in Fig. 4. Because of the $\mathbb{Z}_{2}$-symmetry about the $x=0$ axis in Fig. 4, the junction $I$ is expected to be on $x=0$ and we denote it by $I(0,y_{I})$. For general two points $(x_{1},y_{1})$ and $(x_{2},y_{2})$ on the Poincaré half plane, the distance between them is
@@
Therefore, the distance between $I$ and the RT surface of $c$ is given by
$\min\nolimits_{\phi}D\bigl((0,y_{I}),(-a+r\cos\phi,r\sin\phi)\bigr)\ ,$ (18)
with $(-a+r\cos\phi,r\sin\phi)$ a point on the RT surface of $c$. The distances between $I$ and the RT surface of $a,b$ can be worked out similarly and we finally end up with the length of the bulk dual web
@@
where the first and second term in $\cdots$] comes from the distance between $I$ and the RT surface of $b$ and $\{a,c\}$. The minimization over $y_{I}$ in (19) will be evaluated numerically (see the blue curves in Fig. 5) when comparing with the field-theoretical results.
Field-theoretical calculation. Now we do the calculation in CFT2. Using the replica trick, we first represent the reflected multi-entropy as a path integral on a specific Riemann surface. Here, a new replica number $m$ is introduced for canonical purification (see Ref. [13] for more details about this replica trick).
The $n$-th Rényi version of reflected multi-entropy is
$S_{R}^{(\mathtt{q})}[n]=\frac{1}{1-n}\frac{1}{n^{\mathtt{q}-2}}\log\frac{\mathcal{Z}_{n^{\mathtt{q}-1},m}}{(\mathcal{Z}_{1,m})^{n^{\mathtt{q}-1}}}\ ,$ (20)
and the reflected multi-entropy is obtained by taking the limit $n\to 1$ and $m\to 1$,
$S_{R}^{(\mathtt{q})}=\lim_{m\to 1}\lim_{n\to 1}\frac{1}{1-n}\frac{1}{n^{\mathtt{q}-2}}\log\frac{\mathcal{Z}_{n^{\mathtt{q}-1},m}}{(\mathcal{Z}_{1,m})^{n^{\mathtt{q}-1}}}\ .$ (21)
Now we use the twist operator correlation function to express $\mathcal{Z}_{n^{\mathtt{q}-1},m}$, the path integral on a glued Euclidean spacetime with $mn^{\mathtt{q}-1}$ copies. The conformal weights $h_{i}$ of twist operators $\sigma_{i}(x_{i})$ can be read from the representation of cyclic groups [47],
@@
We leave the derivation of the conformal weights (22) and (23) to Supplemental Material [48]. When $\mathtt{q}=2$, one can check that (22) reduces to the conformal weight of the twist operators for the reflected entropy as we expected  [^5] For reflected entropy, we have [13] $h_{g_{A}}=h_{g_{A}^{-1}}=\frac{cn(m^{2}-1)}{24m}\ ,$ where $g_{A}$ and $g_{A}^{-1}$ are the twist operators located at the two endpoints of interval $A$.. The conformal weights $h_{a}$ of the leading operators $\sigma_{a}$ in operator product expansion $\sigma_{i}(x_{i})\sigma_{j}(x_{j})\to\sigma_{a=ij}(x_{a})$ are
$h_{a=ij}=\frac{c}{12}\left(n^{\mathtt{q}-1}-n^{\mathtt{q}-3}\right)\ ,$ (23)
which is twice the conformal weight of the twist operators for the original multi-entropy as we expected  [^6] The conformal dimension of the multi-entropy twist operator is given in Eq. (29) of [34]. . It can also be checked that for $\mathtt{q}=2$, (23) reduces to twice the conformal weight of the twist operators for EE.
@@
with $\sigma_{i}$ located at $x_{i}$. At large $c$ limit, the calculation of (24) can be converted to solving a monodromy problem [51, 52]. Following the approach in [53] we compute the six-point function (24) at large $c$ limit [48] and the partial derivatives of $S_{R}^{(3)}(A;B;C)$ with respect to $a,b,r$ are plotted in Fig. 5.